calculate the net present value

Assuming monetary benefits of an information system at $85,000 per year, one-time costs of $75,000, recurring costs of $35,000 per year, a discount rate of 12 percent,and a five-year time horizon, calculate the net present value of these costs and benefits of an information system. Also, calculate the overall return oninvestment
Answer #1

Economic feasibility analysis for an Information System (1)

Consider the given data:

Monetary Benefits of IS = $85,000 per year.

One-time costs = $75,000

Recurring costs = $35,000 per year.

Discount rate = 12%.

Time Period = 5 years.

The worksheet given below (figure) shows the Present Value calculations of all costs and benefits, Break-Even Analysis, Overall Return on Investment for this problem using the above data.

The spreadsheet displays the following items:

1. Net Present Values (NPV) of all Benefits and Costs

2. Overall Return On Investment (ROI)

3. Break-Even Analysis (BEA)

________________________________________________________________________

1. Net present value of Benefits and Costs:

Present Value of Benefits or Costs can be calculated using the below formula:

Here, PVn is the present value of Y dollars n years from now, and i is the discount rate.

Present Value (PV) calculations for Benefits:

Benefits start from year 1, so the calculation of PV from year 1 onwards.

Net Present Value of Benefits: The net value of benefit will be the sum of overall benefits and will be calculated using following formulae:

(Cell H8)

Present Value (PV) calculations for Costs: Here, the one-time cost ($75,000) is treated as cost occurring in year 0 (now).

Recurring cost ($35,000) happens every year starting at year 1.

Net Present Value of Costs,

(Cell H16)

________________________________________________________________________

2. Overall Return On Investment (ROI)

From the spreadsheet, Overall NPV is 105,240 (Cell H18) and

NPV of All COSTS is 201,168 (Cell H16).

Therefore, the overall

(Cell H20)

3. Break-Even Analysis (BEA)

This analysis is carried out by first determining the NPV of cash flows on yearly basis.

Subtract the present value of recurring cost from the yearly benefits for computing the Yearly NPV Cash Flows. The Overall NPV Cash Flow is the total flow of cash of all the preceding years.

If line 24 of spreadsheet is observed, it would be find that occurrence of break-even happens between year one and two. This is because year two is the very first year in which Overall NPV is non-negative. (Positive)

Using the info from below Figure,

Therefore, Project break-even occurs at approximately years (rounded).

6PE
Answer #2
Assume that the 85k benefit and 35k cost both occur at the end of each year. Set up the cash flows and discount by year.

T0 -75k
t1 (85k -35k) / 1.12 = 44,642
t2 (85k -35k) / 1.12^2 = 39,859
t3 (85k -35k) / 1.12^3 = 35,589
t4 (85k - 35k) / 1.12^4 = 31,775
t5 (85k -35k) / 1.12^5 = 28,371

NPV = $105,236

The project breaks-even after 2 years because the NPV at that point is already positive. If the costs and benefits were throughout the year, you could calculate a fraction of the year, which is a good reason to simplify by assuming year end flows.